# Dynamic Strategy Adaptation in Multi-Agent Environments with Large Language Models

### v2

@@ Abstract
Large language models (LLMs) demonstrate strong reasoning abilities across mathematical, strategic, and linguistic tasks, yet little is known about how well they reason in dynamic, real-time, multi-agent scenarios, such as collaborative environments in which agents continuously adapt to each other’s behavior, as in cooperative gameplay settings. In this paper, we bridge this gap by combining LLM-driven agents with strategic reasoning and real-time adaptation in cooperative, multi-agent environments grounded in game-theoretic principles such as belief consistency and Nash equilibrium. The proposed framework applies broadly to dynamic scenarios in which agents coordinate, communicate, and make decisions in response to continuously changing conditions. We provide real-time strategy refinement and adaptive feedback mechanisms that enable agents to dynamically adjust policies based on immediate contextual interactions, in contrast to previous efforts that evaluate LLM capabilities in static or turn-based settings. Empirical results show that our method achieves up to a 26% improvement in return over PPO baselines in high-noise environments, while maintaining real-time latency under 1.05 milliseconds. Our approach improves collaboration efficiency, task completion rates, and flexibility, illustrating that game-theoretic guidance integrated with real-time feedback enhances LLM performance, ultimately fostering more resilient and flexible strategic multi-agent systems.
1 Introduction
@@
Large language models (LLMs) are an effective machine learning tool that can perform a broad range of reasoning tasks, such as mathematical problem solving, code generation, commonsense inference, and strategic board games. However, their ability to operate effectively in real-time, multi-agent coordination settings remains under-explored [1]. Given the rise of multi-agent systems, characterized by dynamic collaboration and automation, there is a need for new frameworks capable of supporting real-time strategic adaptation. Although game theory offers a robust mathematical framework for modeling strategic interactions, it typically assumes fixed decision rules that lack the flexibility to accommodate real-world complexity and change [1]. In contrast, reinforcement learning (RL) provides a flexible mechanism for learning through experience; however, it requires extensive training and often produces strategies that lack clear interpretability. To guide this investigation, we ask the following research question: How can LLMs dynamically adapt and refine their game strategies based on real-time feedback in multi-agent, multi-turn environments to achieve enhanced rationality and equilibrium outcomes?
A number of prior works have [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] studied negotiation, tool integration, real-time planning, and embodied reasoning in the context of LLM-based decision-making. For example, He et al. [2] and Lewis et al. [3] explore end-to-end learning for negotiation dialogues, while Toolformer [4] demonstrates how LLMs can autonomously learn to invoke external tools during generation. PaLM-E [5] highlights the value of integrating visual and proprioceptive inputs for grounded multimodal reasoning. ReAct prompting [6] illustrates how models can reason and act by interleaving tool calls with intermediate steps. NegotiationArena [7] and StarCraft II benchmarks [8] provide multi-agent coordination evaluations, and CICERO [9] showcases strategic interaction through dialogue in high-stakes, turn-based environments. Recent frameworks such as CAMEL [10], AutoGen [11], LangAgent [12], and Voyager [13] have introduced collaborative LLM teams with structured roles and planning objectives.
@@
Other solutions such as SWE-agent [14] and MetaGPT [15] propose programming-based role assignment and skill composition pipelines, yet often operate with fixed dialogue states and precomputed coordination graphs. OctoAgents [16] explore zero-shot LLM team formation but require architectural homogeneity and lack continual learning mechanisms. The MindAct [17] and CoLLM [18] models examine collaboration over temporal chains but still assume a known world model. While prior solutions [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] enable symbolic reasoning, tool usage, and basic interagent dialogue, they typically do not support continuous strategic adaptation in cooperative, real-time environments. Most of these existing approaches operate in static settings or rely on single-shot policy inference. There is little ability for agents to revise goals, respond to degraded teammate performance, or restructure plans in response to unforeseen environment transitions. A central open challenge is, thus, how to enable agents to adjust their strategy mid-execution using context-aware feedback, memory persistence, symbolic priors, and feedback-aware control loops that reflect real-time game-theoretic dynamics.
An examination of the literature that addresses uncertainty, partial observability, and strategic adaptation in games reveals several key limitations [19, 20, 21, 13, 22, 23, 24, 9, 25]. GTBench [19] tests LLMs on game-theoretic matrix tasks, finding poor robustness in partially observable environments. GameTheoryLLM [20] formalizes Bayesian agent workflows for negotiation but imposes synchronous update protocols. RationalLLM [21] studies equilibrium behavior but does not support belief evolution or policy trajectory tracking. MARL benchmarks like MAgent [26] and EnvAgent [27] test emergent roles but fail to align agents with structured symbolic constraints. Systems like MindAgent [28], DEAM [29], and ToM-GPT [30] simulate theory-of-mind and belief graphs but rely on heavily curated scenarios. CoaLLM [31] and OpenAgents [32] attempt to unify self-play and LLM planning, though they still assume pre-scripted task flows.
Although CICERO [9] offers a rich symbolic policy in the context of Diplomacy, it is limited to turn-based play. Reinforcement learning from human feedback (RLHF) [25] allows reward shaping but lacks dynamic revision based on environment evolution. Existing frameworks such as like RolePlay [33] and ThoughtAgents [34] aim to inject personality and dialogue abstraction but do not connect memory or belief state with performance feedback. Few systems track and utilize contextual priors across time steps. Even fewer are capable of aligning their behavioral revisions with task-level metrics like regret, coordination efficiency, or Nash deviation. We argue that bridging this gap requires dynamic feedback systems that are tightly coupled with symbolic modeling and game-theoretic planning principles. The field still lacks a unified method for enabling LLM-based agents to revise their strategies in real-time as goals shift, agents deviate, or coordination structures deteriorate.
@@
The strategic use of LLMs in dynamic, real-time, multi-agent environments faces a significant open problem: how to enable agents to reason adaptively, coordinate under uncertainty, and revise decisions as environments evolve. While early systems [19, 20, 21, 13, 22, 23, 9] have made progress in specific areas, such as static equilibrium modeling, symbolic dialogue protocols, or role-based planning, they remain limited in three key aspects: 1) they are unable to revise beliefs in response to shifting goals or noisy conditions, 2) they lack real-time mechanisms for coordination across agents, and 3) they do not incorporate symbolic priors or game-theoretic principles into feedback loops. These limitations highlight the need for integrated multi-agent reasoning frameworks that support real-time strategy adaptation and structured coordination, guided by symbolic feedback and contextual understanding under uncertainty.
@@
The main contribution of this paper is to address the limitations of static decision-making methods, including rule-based strategies, turn-based reasoning, and offline reinforcement learning, by proposing a new framework that enables agents to adapt strategies in real time using LLM-guided reward shaping. By integrating symbolic evaluations from a frozen language model with a reinforcement learning loop, the proposed system refines agent behavior based on contextual signals such as order progress and timestep. This hybrid framework is designed to support collaboration in unpredictable, noisy environments where conventional policies tend to fail. In addition to developing the architecture, the proposed framework is empirically evaluated across dynamic environments to assess performance, robustness to noise, and latency tradeoffs. The proposed method delivers strategic benefits while preserving real-time inference speeds. In summary, the key contributions of this paper are:
@@
  • In terms of real-time performance, PPO-LLM agents operated with an average per-action latency of 1.03 to 1.05 milliseconds, inclusive of prompt generation and LLM inference. This satisfies Overcooked-AI’s latency thresholds and confirms the framework’s suitability for deployment in real-time multi-agent infrastructures with frequent decision cycles.
  • PPO-LLM agents also demonstrated a 35% higher full-task completion rate under noise, reflecting more reliable long-horizon coordination and better adaptation to mid-task disruptions. This includes recovery from failed pickups, bottlenecked transitions, and role misassignments under noisy state transitions.
  • The proposed approach allowed the agents to achieve a 22.7% reduction in average idle steps across evaluation episodes, indicating smoother temporal coordination, less waiting, and tighter synchronization between role-dependent sub-tasks. This reflects better implicit communication and increased responsiveness to evolving task conditions.
@@
The rest of this paper is organized as follows. Section 2 presents the system architecture, including the real-time shaping pipeline, symbolic feedback loop, and PPO policy integration. Section 3 provides empirical results from Overcooked-AI environments, including performance under noise, task completion rate, latency tradeoffs, and convergence to equilibrium. Section 4 concludes with a summary of findings, limitations, and directions for future research on real-time strategic adaptation using language-based feedback.
2 System Architecture
Figure 1: Real-Time Shaping System with PPO + Frozen LLM. The environment emits a grid-state at each timestep. That grid is flattened into a 1D observation vector and passed to both the PPO agent and a prompt generator. In parallel, a prompt generator encodes task context (e.g., remaining orders and step count) into a text prompt. This prompt is evaluated by a frozen LLM (GPT-Neo 1.3B), whose binary "good"/"bad" signal is mapped to a shaping bonus via the reward-shaping module. The PPO agent then updates its policy using this shaped reward signal.
To support dynamic strategy adaptation in cooperative multi-agent environments, we introduce a real-time reward shaping architecture that integrates PPO with feedback from a frozen LLM. This system is designed to enhance agent coordination through symbolic reasoning and structured evaluations, without modifying the LLM weights or retraining during execution. It consists of six interconnected modules that operate simultaneously during each timestep: i) environment state emitter, ii) observation preprocessor, iii) PPO policy network, iv) prompt generation module, v) frozen LLM evaluator, and vi) reward shaping unit, as shown in Figure 1.
At time $t$, the environment emits a grid state $s_{t}$, which represents the spatial layout of $N=2$ agents, interactive objects (such as ingredients and plates), and task-relevant regions (such as delivery counters). Each grid cell encodes the type and state of any entity occupying it. This state is transformed into a one-dimensional observation vector to ensure compatibility with both the PPO policy network and the LLM prompt generator. We instantiate this structure using the Overcooked-AI framework [35], a cooperative multi-agent simulator in which agents must coordinate to cook and deliver meals under time pressure and spatial constraints. Figure 2 illustrates our specific “Burger Kitchen” configuration, where one agent prepares cooked patties while the other assembles and delivers finished burgers. This layout introduces strict temporal dependencies and spatial bottlenecks that require dynamic role allocation and real-time strategic coordination.
The PPO agent, implemented using Stable-Baselines3 [36], takes the observation and computes an action $a_{t}$ based on its learned policy $\pi_{\theta}(a_{t}\mid s_{t})$, where $\theta$ represents the parameters of the neural network. The policy maps observed features to a discrete set of agent behaviors, such as movement or object interaction.
@@ 2 System Architecture
Here, $\gamma$ is the discount factor, and $V(s)$ is the learned value function for state $s$.
@@
Because the LLM remains frozen, its evaluations are deterministic and repeatable. These evaluations instantiate two fundamental game-theoretic principles. The first is belief revision, which parallels Bayesian updating: the LLM adjusts its cooperative judgment in response to evolving prompt content that reflects the current environment state, such as the number of pending orders, role distribution, or elapsed time. This mechanism allows the agent to revise expectations mid-task without retraining. The second principle is strategic consistency, which is implemented by reinforcing only those actions that align with collaborative success. Agents receive positive shaping signals only when their behavior contributes meaningfully to shared objectives, thereby discouraging selfish or unstable strategies and encouraging long-term alignment.
To operationalize these principles, we embed structured symbolic evaluation into the reinforcement learning loop. A prompt generation module continuously encodes task-relevant context, such as pending order count, current timestep, and relative agent positions, which is passed to a frozen LLM that produces a binary cooperative judgment. This symbolic signal is then mapped to a scalar bonus and integrated into each agent’s PPO training pipeline in real time. By providing feedback on the overall quality of collaboration, rather than low-level action success, the framework guides agents toward strategic behaviors that prioritize task flow, anticipation of partner actions, and recovery from errors. This allows for dynamic behavioral adjustments during execution, even in the presence of degraded perception or unpredictable task disruptions. As a result, agents adapt their policies mid-episode using high-level, interpretable feedback. This architecture bridges low-level reinforcement learning with symbolic reasoning, enabling robust coordination, adaptive role assignment, and resilience to uncertainty, including observation noise, partial observability, and real-time variability.
@@
Figure 2: Custom Overcooked-AI “Burger Kitchen” scenario used in our experiments. Agent A (blue) retrieves raw patties from the fridge and cooks them on the Stove, while Agent B (green) assembles cooked patties with buns and condiments at the Burger Station and delivers the finished burger through the Delivery Window. Dashed arrows indicate task flow between stations. This scenario is a good illustration of our model because it imposes strict temporal dependencies and spatial constraints that require real-time role specialization, dynamic coordination, and mid-episode strategy adjustment. These are exactly the types of challenges our PPO+LLM shaping framework is designed to address.
3 Simulation Results and Analysis
We evaluate our framework in the Overcooked-AI simulator[35], a cooperative multi-agent benchmark in which two agents must complete cooking tasks under time pressure and partial observability. In our custom burger kitchen scenario (Figure 2), Agent A must wait while the patty cooks, then hand it off to Agent B for assembly and delivery. This layout enforces role specialization, timing dependencies, and dynamic coordination challenges that test the ability of our PPO+LLM system to adapt strategies based on real-time symbolic feedback.
Figure 3: Mean return across clean and visibility noise environments. PPO+LLM performs comparably in clean settings and maintains advantage under perceptual uncertainty.
Figure 4: Mean return across noisy environments (Order Timing Noise and Combined Noise). PPO+LLM demonstrates improved robustness over PPO under multi-agent coordination disruptions.

### v3

@@ Abstract
Large language models (LLMs) demonstrate strong reasoning abilities across mathematical, strategic, and linguistic tasks. However, little is known about how well they reason in dynamic, real-time, multi-agent scenarios, such as collaborative environments where agents continuously adapt to each other’s behavior, as in cooperative gameplay settings. In this paper, we bridge this gap by introducing a novel framework that integrates LLM-based symbolic evaluation with reinforcement learning for real-time adaptation in cooperative, multi-agent environments without modifying the LLM weights or retraining during execution. Grounded in game-theoretic principles such as belief consistency and equilibrium stability, the framework enables agents to refine strategies during execution without retraining or centralized coordination. Our approach supports dynamic scenarios in which agents coordinate, communicate, and make decisions under continuously changing conditions. We provide real-time strategy refinement and adaptive feedback mechanisms that guide agents toward stable joint policies resembling Nash equilibria, achieved empirically through iterative feedback and symbolic reward shaping. In contrast to previous efforts that evaluate LLM capabilities in static or turn-based settings, our method supports continuous mid-episode revision. Empirical results show that our framework achieves up to a 26% improvement in return over Proximal Policy Optimization baselines in high-noise environments, while maintaining real-time latency under 1.05 milliseconds. These findings demonstrate that integrating LLM-based symbolic evaluations with reinforcement learning enables more resilient, interpretable, and flexible multi-agent systems capable of sustaining robust collaboration in dynamic settings.
Index Terms:
Game Theory, Large Language Models, Multi-Agent Reinforcement Learning, Real-Time Strategy Adaptation, Reward Shaping, Symbolic Feedback
I Introduction
@@
Large language models (LLMs) are an effective machine learning tool that can perform a broad range of reasoning tasks, such as mathematical problem solving [1], code generation [2, 3], commonsense inference [4, 5], and strategic board games [6, 7]. However, their ability to operate effectively in real-time, multi-agent coordination settings remains under-explored [8]. Given the rise of multi-agent systems, characterized by dynamic collaboration and automation, there is a need for new frameworks capable of supporting real-time strategic adaptation. Although game theory offers a robust mathematical framework for modeling strategic interactions, it typically assumes fixed decision rules that lack the flexibility to accommodate real-world complexity and change [8]. In contrast, reinforcement learning (RL) provides a flexible mechanism for learning through experience. However, it requires extensive training and often produces strategies that lack clear interpretability. To guide this investigation, we ask the following research question: How can LLMs dynamically adapt and refine their game strategies based on real-time feedback in multi-agent, multi-turn environments to achieve enhanced rationality and equilibrium outcomes?
@@
A number of prior works have studied negotiation, tool integration, real-time planning, and embodied reasoning in the context of LLM-based decision-making [9, 10, 11, 12, 13, 14, 15, 6, 16, 17, 18, 19]. For example, He et al. [9] and Lewis et al. [10] explore end-to-end learning through LLMs for negotiation dialogues, while Toolformer [11] demonstrates how LLMs can autonomously learn to invoke external tools during generation. PaLM-E [12] highlights the value of integrating visual and proprioceptive inputs for grounded multimodal reasoning. ReAct prompting [13] illustrates how models can reason and act by interleaving tool calls with intermediate steps. NegotiationArena [14] and StarCraft II benchmarks [15] provide multi-agent coordination evaluations, and CICERO [6] showcases strategic interaction through dialogue in high-stakes, turn-based environments. Recent frameworks such as CAMEL [16], AutoGen [17], LangAgent [18], and Voyager [19] have introduced collaborative LLM teams with structured roles and planning objectives. Other solutions such as SWE-agent [2] and MetaGPT [3] propose programming-based role assignment and skill composition pipelines, yet often operate with fixed dialogue states and precomputed coordination graphs. OctoAgents [20] explore zero-shot LLM team formation but require architectural homogeneity and lack continual learning mechanisms. The MindAct [21] and CoLLM [22] models examine collaboration over temporal chains but still assume a known world model. While prior solutions [9, 10, 11, 12, 13, 14, 15, 6, 16, 17, 18, 19] enable symbolic reasoning, tool usage, and basic interagent dialogue, they typically do not support continuous strategic adaptation in cooperative, real-time environments. However, most existing approaches operate in static settings or rely on single-shot policy inference [7, 23, 24, 11, 13]. There is little ability for agents to revise goals, respond to degraded teammate performance, or restructure plans in response to unforeseen environment transitions. A central open challenge is, thus, how to enable agents to adjust their strategy mid-execution using context-aware feedback, memory persistence, symbolic priors, and feedback-aware control loops that reflect real-time game-theoretic dynamics.
@@
The strategic use of LLMs in dynamic multi-agent environments face significant challenges in enabling agents to reason adaptively, coordinate under uncertainty, and revise decisions as environments evolve. While early systems [7, 23, 24, 19, 25, 26, 6] have made progress in specific areas, such as static equilibrium modeling, symbolic dialogue protocols, or role-based planning, they remain limited in three key aspects: 1) they cannot revise their beliefs which means they cannot update internal models of the environment or other agents’ intentions when goals shift or conditions become noisy, 2) they lack real-time mechanisms for coordination across agents, and 3) they do not incorporate symbolic priors (structured knowledge constraints) or game-theoretic principles (strategic reasoning frameworks such as equilibrium or bargaining models) into feedback loops. These limitations highlight the need for integrated multi-agent reasoning frameworks that support real-time strategy adaptation and structured coordination, guided by symbolic feedback and contextual understanding under uncertainty.
Although CICERO [6] offers a rich symbolic policy in the context of Diplomacy, it is limited to turn-based play. Reinforcement learning from human feedback (RLHF) [27] allows reward shaping but lacks dynamic revision based on environment evolution. Existing frameworks such as RolePlay [28] and ThoughtAgents [5] aim to inject personality and dialogue abstraction but do not connect memory or belief state with performance feedback. Few systems track and utilize contextual priors across time steps. Even fewer are capable of aligning their behavioral revisions with task-level metrics like regret, coordination efficiency, or Nash deviation. We argue that bridging this gap requires dynamic feedback systems that are tightly coupled with symbolic modeling and game-theoretic planning principles. This naturally brings us back to the broader challenge of continuous strategic adaptation in cooperative, real-time environments.
Building on this motivation, we examine continuous strategic adaptation more directly. The key difficulty is enabling LLM-based agents to revise their strategies in real-time as goals shift, agents deviate, or coordination structures deteriorate. While prior work introduces symbolic reasoning or dialogue abstractions, these systems fall short of coupling behavioral revisions with long-horizon planning, leaving strategic adaptation incomplete in evolving environments.
Beyond continuous adaptation, agents must also cope with uncertainty, incomplete information, and evolving opponent behavior. In practice, they must not only adapt their strategies continuously, but also reason under partial observability and dynamic conditions. An examination of the literature that addresses uncertainty, partial observability, and strategic adaptation in games reveals several key limitations [7, 23, 24, 19, 25, 26, 6]. GTBench [7] tests LLMs on game-theoretic matrix tasks, finding poor robustness in partially observable environments. GameTheoryLLM [23] formalizes Bayesian agent workflows for negotiation but imposes synchronous update protocols. RationalLLM [24] studies equilibrium behavior but does not support belief evolution or policy trajectory tracking. Multi-Agent Reinforcement Learning benchmarks like MAgent [29] and EnvAgent [30] test emergent roles but fail to align agents with structured symbolic constraints. Systems like MindAgent [4], DEAM [31], and ToM-GPT [32] simulate theory-of-mind and belief graphs but rely on heavily curated scenarios. CoaLLM [33] and OpenAgents [34] attempt to unify self-play and LLM planning, though they still assume pre-scripted task flows. Taken together, these approaches demonstrate progress in modeling interaction and adaptation, yet they remain fragmented and constrained by narrow experimental setups. None provide a unified framework for handling uncertainty, partial observability, and continuous adaptation in open-ended, real-time environments.
@@
The main contribution of this paper is to address the limitations of static decision-making methods, including rule-based strategies, turn-based reasoning, and offline reinforcement learning, by proposing a novel framework that enables agents to adapt strategies in real time using LLM-guided reward shaping. By integrating symbolic evaluations from a frozen language model with a reinforcement learning loop, the proposed system refines agent behavior based on contextual signals such as order progress (the evolving state of tasks or objectives being completed within the environment) and timestep.This hybrid framework is designed to support collaboration in unpredictable, noisy environments where conventional policies tend to fail. In addition to developing the architecture, the proposed framework is empirically evaluated across dynamic environments to assess performance, robustness to noise, and latency tradeoffs. The proposed method delivers strategic benefits while preserving real-time inference speeds. In summary, the key contributions of this paper are:
@@
  • PPO-LLM agents achieved real-time suitability with per-action latencies of 1.03–1.05 milliseconds while also demonstrating stronger coordination: a 35% higher full-task completion rate under noise and a 22.7% reduction in idle steps across episodes. Together, these results indicate smoother temporal synchronization, improved recovery from disruptions, and more reliable long-horizon performance compared to PPO baselines.
@@
The rest of this paper is organized as follows. Section II presents the system architecture, including the real-time shaping pipeline, symbolic feedback loop, and PPO policy integration. Section III provides empirical results from Overcooked-AI environments, including performance under noise, task completion rate, latency tradeoffs, and convergence to equilibrium. Section IV concludes with a summary of findings, limitations, and directions for future research on real-time strategic adaptation using language-based feedback.
II System Architecture
Figure 1: Real-Time Shaping System with PPO + Frozen LLM. The environment emits a grid-state at each timestep. That grid is flattened into a 1D observation vector and passed to both the PPO agent and a prompt generator. In parallel, a prompt generator encodes task context (e.g., remaining orders and step count) into a text prompt. This prompt is evaluated by a frozen LLM (GPT-Neo 1.3B), whose binary “good”/“bad” signal is mapped to a shaping bonus via the reward-shaping module. The PPO agent then updates its policy using this shaped reward signal.
We instantiate our experimental environment using Overcooked-AI [35], a cooperative multi-agent simulator in which agents must coordinate to cook and deliver meals under time pressure and spatial constraints. At each time step $t$, the environment emits a grid state $s_{t}$ that captures the spatial layout of $N=2$ agents, interactive objects (such as ingredients and plates), and task-relevant regions (such as delivery counters). Each grid cell encodes the type and state of any entity occupying it, and the grid is subsequently transformed into a one-dimensional observation vector to ensure compatibility with both the PPO policy network and the LLM prompt generator. Figure 2 illustrates our specific “Burger Kitchen” configuration, where one agent prepares cooked patties while the other assembles and delivers finished burgers. This layout introduces strict temporal dependencies and spatial bottlenecks that require dynamic role allocation and real-time strategic coordination.
The PPO agent, implemented using Stable-Baselines3 [36], takes the observation $o_{t}$ and computes an action $a_{t}$ based on its learned policy $\pi_{\theta}(a_{t}\mid o_{t})$, where $\theta$ represents the parameters of the neural network. The policy maps observed features to a discrete set of agent behaviors, such as movement or object interaction.
@@ II System Architecture
Using these temporal difference errors, the GAE is computed as
@@
Because the LLM remains frozen, its evaluations are deterministic and repeatable. In our implementation, we directly query the model logits for specific tokens (e.g., “good” vs. “bad”) without any sampling, ensuring outputs are stable and reproducible across identical prompts. To maintain stability during training, the shaping coefficient $\lambda$ is kept small and explicitly bounded, preventing destabilization or runaway updates. Formally, if the shaped reward is defined as
$r^{\prime}_{t}=r_{t}+\lambda s_{t},\quad s_{t}\in[-1,1].$ (5)
then the perturbation per step satisfies $|r^{\prime}_{t}-r_{t}|\leq\lambda$. Over the discounted horizon, and for a fixed policy $\pi$, the cumulative shift is bounded by
$|V^{\prime\pi}(s)-V^{\pi}(s)|\leq\frac{\lambda}{1-\gamma}.$ (6)
The corresponding shift in advantages is
$|A^{\prime\pi}(s,a)-A^{\pi}(s,a)|\leq\frac{2\lambda}{1-\gamma}.$ (7)
These bounds ensure that both value targets and advantages remain close to their original scale during PPO training.
Beyond stability, these evaluations instantiate two fundamental game-theoretic principles. The first is belief revision, which parallels Bayesian updating. Let $b_{t}$ denote the LLM’s belief score of cooperative success given prompt $\phi(s_{t})$. We model this as
$b_{t+1}=\frac{P(\phi(s_{t+1})\mid\mathrm{coop})\,b_{t}}{P(\phi(s_{t+1})\mid\mathrm{coop})\,b_{t}+P(\phi(s_{t+1})\mid\lnot\mathrm{coop})\,(1-b_{t})}.$ (8)
so that the LLM revises its cooperative judgment in response to evolving prompt content such as pending orders, role distribution, or elapsed time. This allows the agent to update expectations mid-task without retraining, while still remaining deterministic.
The second principle is strategic consistency, which is implemented by reinforcing only those actions that align with collaborative success. Formally, the shaping signal is applied as
$s_{t}=\begin{cases}+1&\text{if }a_{t}\in\mathcal{A}_{\text{coop}},\\ 0&\text{otherwise}.\end{cases}$ (9)
where $\mathcal{A}_{\text{coop}}$ is the set of actions contributing meaningfully to shared objectives. This ensures agents receive positive shaping signals only when their behavior advances cooperative goals, thereby discouraging selfish or unstable strategies and promoting long-term alignment.
Finally, to guarantee that shaping never dominates the task reward, $\lambda$ must satisfy
$\lambda\leq\min\!\left\{(1-\gamma)\,\delta_{V},\;\tfrac{1-\gamma}{2}\,\varepsilon_{\text{clip}}\,A_{\text{scale}}\right\}.$ (10)
where $\delta_{V}$ is the maximum tolerated critic-target drift, $\varepsilon_{\text{clip}}$ is the PPO clipping parameter, and $A_{\text{scale}}$ is a characteristic scale of the unshaped advantages. This bound ensures that shaping acts only as a small regularizer rather than overwhelming the base reward signal.
To operationalize these principles, we embed structured symbolic evaluation into the reinforcement learning loop through the LLM. A prompt generation module continuously encodes task-relevant context, such as the number of pending orders, the current timestep, and relative agent positions. This structured description is passed to a frozen LLM, which produces a binary cooperative judgment based on its logits for tokens such as “good” and “bad.” Formally, if the logit for “good” exceeds that for “bad,” a scalar bonus $\lambda$ is applied as in Eq. (1), yielding a shaped reward $r^{\prime}_{t}=r_{t}+\lambda$ (Eq. LABEL:eq:shaped_reward_simple). Otherwise, the base reward $r_{t}$ is retained. This ensures that feedback reflects the overall quality of collaboration rather than isolated action outcomes.
The shaped reward $r^{\prime}_{t}$ is then used to compute temporal-difference errors $\delta_{t}$ (Eq. 3) and advantages $\hat{A}_{t}$ via generalized advantage estimation (Eq. 4), allowing cooperative feedback to influence long-horizon credit assignment in PPO. By integrating symbolic evaluations in this way, the framework guides agents toward strategic behaviors that emphasize task flow, anticipation of partner actions, and recovery from errors.
To ensure stability, we bound the effect of shaping at every stage. Equation 5 guarantees that per-step perturbations remain within $|r^{\prime}_{t}-r_{t}|\leq\lambda$, while Eqs. 6 and 7 show that the cumulative shifts in value and advantage estimates are capped. These guarantees prevent destabilization during PPO training even under noisy or partial observations.
Beyond stability, the LLM feedback instantiates two game-theoretic principles. First, belief revision (Eq. 8) parallels Bayesian updating: the LLM revises its cooperative judgment deterministically as the prompt evolves with state changes such as order backlog or elapsed time. Second, strategic consistency (Eq. 9) enforces that shaping signals are positive only when actions belong to the cooperative set $\mathcal{A}_{\text{coop}}$, thereby discouraging selfish or unstable strategies. Finally, Eq. 10 constrains $\lambda$ so that shaping acts only as a regularizer, never overwhelming the task reward.
As a result, agents dynamically adapt their policies mid-episode using high-level, interpretable feedback. This architecture bridges low-level reinforcement learning with symbolic reasoning, enabling robust coordination, adaptive role assignment, and resilience to uncertainty, including observation noise, partial observability, and real-time variability.
@@
Figure 2: Custom Overcooked-AI “Burger Kitchen” scenario used in our experiments. Agent A (blue) retrieves raw patties from the fridge and cooks them on the Stove, while Agent B (green) assembles cooked patties with buns and condiments at the Burger Station and delivers the finished burger through the Delivery Window. Dashed arrows indicate task flow between stations. This scenario imposes strict temporal dependencies and spatial constraints that require real-time role specialization, dynamic coordination, and mid-episode strategy adjustment,exactly the challenges our PPO+LLM shaping framework is designed to address.
III Simulation Results and Analysis
Figure 3: Mean return across clean and visibility-noise environments. PPO+LLM performs comparably in clean settings and maintains advantage under perceptual uncertainty.
Figure 4: Mean return across noisy environments (Order Timing Noise and Combined Noise). PPO+LLM demonstrates improved robustness over PPO under multi-agent coordination disruptions. Results averaged over n=200 evaluation episodes.